# Qini Curve Estimation Under Clustered Network Interference

### v1

@@ 5 Estimation strategies
In this section we consider multiple strategies for Qini curve estimation. In particular, we will focus on weighting estimators that use the propensity score. Since we assumed the propensity score to be known in our design, we avoid the need to model any other nuisance parameters. Throughout this section, due to the similarity of $\phi(\pi)$ and $\psi(\pi)$, we only present strategies for estimating $\phi(\pi)$ which analogously can be applied for estimating $\psi(\pi)$ as well.
@@ 5.1 Cluster-level inverse probability weighting
$\widehat{\phi}^{\text{\;IPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}Y_{i}$
where $\mathbf{1}(\cdot)$ denotes the indicator function. This estimator is a natural extension of the traditional inverse probability weighting (IPW) estimator (Robins et al., 1994) to settings with clustered network interference, see e.g.  Tchetgen and VanderWeele (2012). For this reason, we will refer to $\widehat{\phi}^{\text{\;IPW}}(\pi)$ as the standard IPW estimator. Although we can show that $\widehat{\phi}^{\text{\;IPW}}(\pi)$ is an unbiased estimator under assumption 3.1, its efficiency is poor which becomes evident from inspecting its sampling variance,
$\mathrm{Var}(\widehat{\phi}^{\text{\;IPW}}(\pi))$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\left\{\mathbb{E}\left[\omega(X_{i})\left[Y_{i}(\boldsymbol{\pi}_{i})\right]^{2}\right]+\mathrm{Var}\left(Y_{i}(\boldsymbol{\pi}_{i})\right)\right\}$
$\omega(X_{i})$ $=\left[\prod_{j=1}^{M_{i}}\left(\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right)-1\right]$
We derive the unbiasedness and variance of the standard IPW estimator in Appendix A.2.
Since $e_{1}(X_{i})\neq 0$ and $e_{0}(X_{i})\neq 0$ due to assumption 3.1, the factor $\omega(X_{i})$ increases exponentially with the cluster size $M_{i}$. Consequently, its variance scales exponentially with the cluster size $M_{i}$ which makes it prohibitively difficult to use the standard IPW estimator for Qini curve estimation in scenarios where the cluster size $M_{i}$ is large.
@@ 5.2 Interference under a fractional exposure mapping
One strategy to deal with interference is by defining exposure mappings (Aronow and Samii, 2017). An exposure mapping is a function $d_{ij}:\{0,1\}^{M_{i}}\rightarrow\mathcal{D}$ between all possible treatment configurations for unit $(i,j)$ and a representation (or, embedding) of the treatment configurations. In essence, we want to map similar treatment configurations to the same “effective treatment” (Manski, 2013). If the space $\mathcal{D}$ has smaller cardinality than the original space $\{0,1\}^{M_{i}}$, which has cardinality $2^{M_{i}}$, we have the possibility for more efficient estimation. Any exposure mapping, however, must fulfill the following condition.
Assumption 5.1.
@@
$\widehat{\phi}^{\;\text{fracIPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M_{i}}\frac{\mathbf{1}\left(W_{ij}=\pi_{ij},\overline{W}_{i}=\overline{\pi}_{i}\right)}{q_{ij}(\boldsymbol{\pi}_{i},X_{i})}Y_{ij}$
where $q_{ij}(\boldsymbol{\pi}_{i},X_{i})=\Pr(W_{ij}=\pi_{ij},\overline{W}_{i}=\overline{\pi}_{i}\mid X_{i})$. As we show in Appendix A.3, the probability $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$ can be expressed in terms of the known propensity score.
The fractional IPW estimator $\widehat{\phi}^{\;\text{fracIPW}}$ is an unbiased estimator for the policy value $V(\pi)$ under assumptions 3.1 and 5.1 (see Appendix A.4 for the proof). Compared to the standard IPW estimator, its variance scales more favorably with the cluster size $M_{i}$ because the fractional exposure mapping reduces the cardinality of the treatment space from $2^{M_{i}}$ to $2(M_{i}+1)$ per cluster. This reduction makes estimation more feasible in settings with large clusters.
However, this does not imply that the variance of $\widehat{\phi}^{\;\text{fracIPW}}$ grows linearly with $M_{i}$. The estimator remains inversely proportional to the probability $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$. As $M_{i}$ increases, the number of possible treatment fractions grows, making it less likely to observe a specific fraction. Consequently, $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$ approaches zero as $M_{i}$ increases, which amplifies the variance, though at a slower rate than the standard IPW estimator.
@@ 5.3 Interference under $\beta$-additive model
Next, we consider another strategy that can reduce variance compared to the standard IPW estimator. Specifically, we use the following polynomial model to describe the interference, as proposed by Zhang and Imai (2023).
Assumption 5.2.
@@
This assumption states that each unit’s conditional mean potential outcome is a linear function of the augmented treatment vector $\gamma(\mathbf{W}_{i})$, which includes interaction terms up to order $\beta$ between treatments within the same cluster. For this reason, we refer to the above assumption as the $\beta$-additive assumption.
@@
$\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta)=\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{\mathcal{U}\in\mathcal{I}_{i}^{\beta}}\prod_{j\in\mathcal{U}}\left(\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{\pi_{ij}}(X_{i})}-1\right)\right]Y_{i}~{}.$
To use the estimator $\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta)$, we must specify $\beta$. When this parameter is chosen to satisfy assumption 5.2 alongside assumption 3.1, Zhang and Imai (2023) proved that $\widehat{\phi}^{\;\beta\text{-IPW}}$ is an unbiased estimator for the policy value $V(\pi)$.
The choice of $\beta$ dictates the strength of the $\beta$-additive assumption, which becomes less restrictive as $\beta$ increases. Setting $\beta=\max_{i}M_{i}$ imposes no additional constraints on the interference structure since, in this case, we have that $\widehat{\phi}^{\;\beta\text{-IPW}}=\widehat{\phi}^{\text{\;IPW}}$ (Zhang and Imai, 2023). Thus, from a practical point of view, the greatest variance reduction can be achieved by using a smaller $\beta$.
To highlight the best variance reduction we can possibly achieve with the $\beta$-IPW estimator, we consider the special case of $\widehat{\phi}^{\text{addIPW}}(\pi)=\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta=1)$, which we refer to as the additive IPW estimator because there are no interactions between multiple treatments within the same cluster. For the additive IPW estimator, we can prove that its variance will scale quadratically with the cluster size, that is, $\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)\propto\max_{i}M_{i}^{2}$ (see Appendix A.5 for proof). This is a notable improvement over the exponential scaling of the standard IPW estimator.
5.4 Selecting the best estimation strategy: considerations from a bias-variance perspective
We have three possible estimators to use: the standard IPW estimator, the fractional IPW estimator, and the $\beta$-IPW estimator. While the standard IPW estimator will suffer from high variance when cluster sizes $M_{i}$ increase, the other two estimators alleviate this issue under specific conditions on the interference. However, this variance reduction may come at the cost of bias if those conditions are violated. Effectively navigating this bias-variance trade-off will be important. Overemphasizing variance reduction may lead to misleading policy evaluations, while prioritizing unbiasedness may produce unstable estimates.
In the next section, we empirically examine which estimation strategy performs best across different settings to provide insights on the bias-variance trade-offs.
@@ 6 Experiments
We aim to evaluate the performance of our proposed strategies for estimating Qini curves under clustered network interference. To balance realistic structures of interference with the benefits of synthetic data, we designed a simulator that mimics a marketplace, common in e-commerce or marketing, where interference arises through cannibalization among product items sold by different vendors. We present our simulator as a framework that can be reused for future research on the topic of interference. We provide all code for reproducing our simulations in the Github repository github.com/bookingcom/uplift-interference-simulator.
In our experiments, we compare five strategies for estimating Qini curves. First, we implement a naive strategy that ignores all interference, which we refer to as the naive estimator (more details are provided in Appendix B). Next, we implement the estimators discussed in this paper: the standard IPW estimator, the fractional IPW estimator and the $\beta$-IPW estimator with $\beta=1$ (additive IPW) or $\beta=2$.
@@
Our experiments focus on two key aspects of using Qini curves for decision-making. The first aspect is calibration: how accurately the estimates $\{\widehat{Q}_{k}\}_{k=0}^{K}$ reflect the ground truth values $\{Q_{k}\}_{k=0}^{K}$. We assess this using bias $K^{-1}\sum_{k=1}^{K}\mathbb{E}[\widehat{Q}_{k}-Q_{k}]$, variance $K^{-1}\sum_{k=1}^{K}\mathrm{Var}(\widehat{Q}_{k})$, and mean squared error $K^{-1}\sum_{k=1}^{K}\mathbb{E}[(\widehat{Q}_{k}-Q_{k})^{2}]$. The second aspect is discrimination: the ability to determine which policy is better. For this, we rank policies based on the estimated area under the Qini curve (higher is better) and use Kendall rank correlation to measure how well each estimator ranks policies compared to the ground truth ranking. By default, we evaluate all methods on a noise-perturbed version of the optimal treatment prioritization rule, which uses oracle knowledge of the data-generating process, to simulate a policy between optimal and random. To simplify evaluation, we perform experiments in the uniform cost case.
@@ 6.1 Simulating an e-commerce marketplace with clustered network interference
In this section, we describe a data-generating process where clusters correspond to potential buyers searching for some item, while treatment units are the items shown. In this marketplace, the treatment of an item could correspond to e.g. discounts or promotions, and the outcome is whether the item was purchased by the buyer. Treatment effects manifest as an incremental change in the probability of a purchase to occur due to the treatment. Each buyer can make at most one purchase, causing cannibalization as treatments may shift purchases between items rather than increasing total purchases.
To construct the dataset, we first sample the covariates and treatment. Next, to introduce heterogeneous treatment effects, we compute an item attractiveness score matrix $\mathbf{A}$, where each element $A_{ij}\in[0,1]$ represents buyer $i$’s interest in purchasing item $(i,j)$. The elements in $\mathbf{A}$ depend on the covariates and assigned treatment. Details on this sampling and computation are provided in Appendix C.1. For simplicity, we assume all buyers observe the same number of items, denoted by $M$.
Next, we sample the outcome $Y_{ij}$, which indicates if item $(i,j)$ is purchased, in two steps. First, we sample the binary outcome $Y_{i}$ for buyer $i$. If a purchase occurs for $i$, we sample which item the buyer purchased. We sample $Y_{i}$ according to the Bernoulli probability $\eta(\mathbf{A}_{i})=P(Y_{i}=1\mid\mathbf{A}_{i})$. The structure of the interference in this dataset is largely determined by the choice of $\eta$.
We consider three alternatives for $\eta$. The simplest is $\eta_{\text{max}}(\mathbf{A}_{i})=\max_{j}A_{ij}$ where only the most attractive item contributes the probability of a purchase by buyer $i$. We refer to $\eta_{\text{max}}$ as the max function. Next, we consider the product function $\eta_{\text{product}}(\mathbf{A}_{i})=1-\prod_{j=1}^{M}(1-A_{ij})$. This function assumes that items contribute independently to a purchase such that $P(Y_{i}=1)=1-P(Y_{i}=0)=1-\prod_{j=1}^{M}P(Y_{ij}=0)$. Lastly, the third function is inspired by position bias, commonly found in ranking systems used in e-commerce platforms (Joachims et al., 2005). We refer to this as the exponential decay function, defined as $\eta_{\text{exp-decay}}(\mathbf{A}_{i})=\sum_{j=1}^{M}(\frac{1}{2})^{\text{rank}(A_{ij})}A_{ij}$, where $\text{rank}(\cdot)$ returns the rank of the attractiveness scores for buyer $i$ in descending order. Each successive item contributes half as much as the preceding one to the probability of a purchase by buyer $i$.
For the final step, if we sample $(Y_{i}=1)$, we determine which item $(i,j)$ is purchased. This is done by sampling according to the probabilities given by the softmax function $P(Y_{ij}=1\mid Y_{i}=1,\mathbf{A}_{i})=\frac{e^{A_{ij}}}{\sum_{j=1}^{M}e^{A_{ij}}}$.

### v2

@@ 4 Estimation strategies
In this section we consider three strategies for estimating Qini curves under clustered network inference. In particular, we will focus on weighting estimators that require the propensity score, $e_{w}(x)=\Pr(W=w\mid X=x)$. Since we assumed the propensity score to be known in our design, we avoid the need to estimate any other nuisance models. Doing so avoids the risk of introducing additional biases due to misspecification when having to estimate nuisance models. Although weighting estimators are the main focus of this paper, we also present in Appendix B some preliminary investigations into techniques for augmenting weighted estimators by incorporating nuisance models that predict the outcome.
Throughout this section, due to the similarity of the estimands for the policy value $\phi(\pi)$ and cost $\psi(\pi)$, we only present strategies for estimating the policy value $\phi(\pi)$ which analogously can be applied for estimating the policy cost $\psi(\pi)$ as well.
@@ 4.1 Cluster-level inverse probability weighting
$\widehat{\phi}^{\text{\;IPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\frac{\mathbf{1}\left(\mathbf{W}_{i}=\boldsymbol{\pi}_{i}\right)}{\prod_{j=1}^{M_{i}}e_{\pi_{ij}}\left(X_{i}\right)}\tilde{Y}_{i}$ (6)
where $\mathbf{1}(\cdot)$ denotes the indicator function. This estimator is a natural extension of the traditional inverse probability weighting (IPW) estimator (Robins et al., 1994) to settings with clustered network interference, see e.g.  Tchetgen & VanderWeele (2012). For this reason, we will refer to $\widehat{\phi}^{\text{\;IPW}}(\pi)$ as the standard IPW estimator. We can show the following important properties of the standard IPW estimator (see proof in Appendix A.2).
Theorem 1.
Under Assumption 1, the standard IPW estimator $\widehat{\phi}^{\text{\;IPW}}(\pi)$ is an unbiased estimator for the policy value, $\mathbb{E}[\widehat{\phi}^{\text{\;IPW}}(\pi)]=V(\pi)$, and its sampling variance can be written as
$\mathrm{Var}\left(\widehat{\phi}^{\text{\;IPW}}(\pi)\right)$ $=\frac{1}{N^{2}}\sum_{i=1}^{N}\left\{\mathbb{E}\left[\omega(X_{i})\left[\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\right]^{2}\right]+\mathrm{Var}\left(\tilde{Y}_{i}(\boldsymbol{\pi}_{i})\right)\right\}$ (7)
$\omega(X_{i})$ $=\left[\prod_{j=1}^{M_{i}}\left(\frac{e_{1}(X_{i})e_{0}(X_{i})}{e_{\pi_{ij}}\left(X_{i}\right)^{2}}+1\right)-1\right]$ (8)
From the above theorem, we observe that while the standard IPW estimator is unbiased, its efficiency is poor which becomes evident from inspecting its sampling variance, Since $e_{1}(X_{i})\neq 0$ and $e_{0}(X_{i})\neq 0$ due to Assumption 1, the factor $\omega(X_{i})$ increases exponentially with the cluster size $M_{i}$. Consequently, its variance scales exponentially with the cluster size $M_{i}$ which makes it prohibitively difficult to use the standard IPW estimator for Qini curve estimation in scenarios where the cluster size $M_{i}$ is large.
@@ 4.2 Interference under a fractional exposure mapping
One strategy to deal with interference is by defining exposure mappings (Aronow & Samii, 2017). An exposure mapping is a function $d_{ij}:\{0,1\}^{M_{i}}\rightarrow\mathcal{D}$ between all possible treatment configurations for unit $(i,j)$ and a representation (or, embedding) of the treatment configurations. In essence, we want to map similar treatment configurations to the same “effective treatment” (Manski, 2013). If the space $\mathcal{D}$ has smaller cardinality than the original space $\{0,1\}^{M_{i}}$, which has cardinality $2^{M_{i}}$, we have the possibility for more efficient estimation. Any exposure mapping, however, must fulfill the following condition.
Assumption 2.
@@
$\widehat{\phi}^{\;\text{fracIPW}}(\pi)=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M_{i}}\frac{\mathbf{1}\left(W_{ij}=\pi_{ij},\overline{W}_{i}=\overline{\pi}_{i}\right)}{q_{ij}(\boldsymbol{\pi}_{i},X_{i})}Y_{ij}$ (9)
where $q_{ij}(\boldsymbol{\pi}_{i},X_{i})=\Pr(W_{ij}=\pi_{ij},\overline{W}_{i}=\overline{\pi}_{i}\mid X_{i})$. As we show in Appendix A.3, the probability $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$ can be expressed in terms of the known propensity score. We can now show that the fractional IPW estimator is unbiased in the setting that the fractional exposure mapping is correctly specified (see proof in Appendix A.4).
Theorem 2.
Under Assumptions 1 and 2, we have that the fractional IPW estimator $\widehat{\phi}^{\;\text{fracIPW}}(\pi)$ is an unbiased estimator for the policy value, $\mathbb{E}\left[\widehat{\phi}^{\;\text{fracIPW}}(\pi)\right]=V(\pi)$.
Notably, the fractional IPW estimator requires additional assumptions compared to the standard IPW estimator to be unbiased, but its variance scales more favorably with the cluster size $M_{i}$ because the fractional exposure mapping reduces the cardinality of the treatment space from $2^{M_{i}}$ to $2(M_{i}+1)$ per cluster. This reduction makes estimation more feasible in settings with large clusters. However, this does not imply that the variance of $\widehat{\phi}^{\;\text{fracIPW}}$ grows linearly with $M_{i}$. The estimator remains inversely proportional to the probability $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$. As $M_{i}$ increases, the number of possible treatment fractions grows, making it less likely to observe a specific fraction. Consequently, $q_{ij}(\boldsymbol{\pi}_{i},X_{i})$ approaches zero as $M_{i}$ increases, which amplifies the variance, though at a slower rate than the standard IPW estimator.
@@ 4.3 Interference under $\beta$-additive model
Next, we consider another strategy that can reduce variance compared to the standard IPW estimator. Specifically, we use the following polynomial model to describe the interference, as proposed by Zhang & Imai (2025).
Assumption 3.
@@
This assumption states that each unit’s conditional mean potential outcome is a linear function of the augmented treatment vector $\gamma(\mathbf{W}_{i})$, which includes interaction terms up to order $\beta$ between treatments within the same cluster. We will refer to the above assumption as the $\beta$-additive assumption.
@@
$\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta)=\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{\mathcal{U}\in\mathcal{I}_{i}^{\beta}}\prod_{j\in\mathcal{U}}\left(\frac{\mathbf{1}(W_{ij}=\pi_{ij})}{e_{\pi_{ij}}(X_{i})}-1\right)\right]\tilde{Y}_{i}~.$ (10)
To use the estimator $\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta)$, we must specify $\beta$. When this parameter is chosen to satisfy Assumption 3 alongside Assumption 1Zhang & Imai (2025) proved the following result.
Theorem 3.
Under Assumptions 1 and 3, we have that the $\beta$-IPW estimator $\widehat{\phi}^{\;\beta\text{-IPW}}(\pi)$ is an unbiased estimator for the policy value, $\mathbb{E}\left[\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta)\right]=V(\pi)$.
The choice of $\beta$ dictates the strength of the $\beta$-additive assumption, which becomes less restrictive as $\beta$ increases. Setting $\beta=\max_{i}M_{i}$ imposes no additional constraints on the interference structure since, in this case, we have that $\widehat{\phi}^{\;\beta\text{-IPW}}=\widehat{\phi}^{\text{\;IPW}}$ (Zhang & Imai, 2025). Thus, from a practical point of view, the greatest variance reduction can be achieved by using a smaller $\beta$.
To highlight the best variance reduction we can possibly achieve with the $\beta$-IPW estimator, we consider the special case of $\widehat{\phi}^{\text{addIPW}}(\pi)=\widehat{\phi}^{\;\beta\text{-IPW}}(\pi;\beta=1)$, which we refer to as the additive IPW estimator because there are no interactions between multiple treatments within the same cluster. Here, we can establish the following claim (see Appendix A.5 for proof).
Lemma 2.
We have that $\mathrm{Var}\left(\widehat{\phi}^{\text{addIPW}}\right)\propto\max_{i}M_{i}^{2}$.
Thus, for the additive IPW estimator, we observe that its variance will scale quadratically with the cluster size. This is a notable improvement over the exponential scaling of the standard IPW estimator.
@@ 5 Experiments
We aim to evaluate the performance of our proposed strategies for estimating Qini curves under clustered network interference. To balance realistic structures of interference with the benefit of having a ground-truth in synthetic data, we designed a simulator that mimics an e-commerce marketplace where interference arises through cannibalization among product items sold by different vendors. Notably, we implemented various interference structures in which Assumptions 2 and 3 are also violated. We present our simulator as a framework that can be reused for future research on the topic of interference. We provide all code in the GitHub repository: https://github.com/bookingcom/uplift-interference-simulator.
In our experiments, we compare five strategies for estimating Qini curves. First, we implement a strategy that ignores all interference, which we refer to as the naive estimator (more details are found in Appendix C). Next, we implement the estimators discussed in this paper: the standard IPW estimator, the fractional IPW estimator and the $\beta$-IPW estimator with $\beta=1$ (additive IPW) or $\beta=2$.
@@
Our experiments focus on two key aspects of using Qini curves for decision-making. The first aspect is calibration: how accurately the estimates $\{\widehat{Q}_{k}\}_{k=0}^{K}$ reflect the ground truth values $\{Q_{k}\}_{k=0}^{K}$. Depending on the experiment, we assess this using bias $K^{-1}\sum_{k=1}^{K}\mathbb{E}[\widehat{Q}_{k}-Q_{k}]$, variance $K^{-1}\sum_{k=1}^{K}\mathrm{Var}(\widehat{Q}_{k})$, and mean squared error $K^{-1}\sum_{k=1}^{K}\mathbb{E}[(\widehat{Q}_{k}-Q_{k})^{2}]$. The second aspect is discrimination: the ability to determine which policy is better. For this, we rank policies based on the estimated area under the Qini curve (higher is better) and use Kendall rank correlation to measure how well each estimator ranks policies compared to the ground truth ranking. By default, since our focus lies on evaluating treatment policies rather than the policies themselves, we implement a simple baseline policy with access to the underlying data-generating process. We then progressively degrade its performance by adding increasing levels of noise to its scoring rule. More details on this policy are provided in Appendix D.1. To simplify evaluation, we perform experiments in the uniform cost case.
@@ 5.1 Simulating an e-commerce marketplace with clustered network interference
In this section, we describe a data-generating process where clusters correspond to potential buyers searching for some item, while treatment units are the items shown. In this marketplace, the treatment of an item $W_{ij}$ could correspond to e.g. discounts or promotions, and the outcome $Y_{ij}$ is whether the item $(i,j)$ was purchased by the buyer $i$. Treatment effects manifest as an incremental change in the probability of a purchase to occur due to the treatment. Each buyer can make at most one purchase, causing cannibalization as treatments may shift purchases between items rather than increasing total purchases.
To construct the dataset, we first sample the covariates $(X_{i},Z_{ij})$ and treatment $W_{ij}$. Next, to introduce heterogeneous treatment effects, we compute an item attractiveness score matrix $\mathbf{A}$, where each element $A_{ij}\in[0,1]$ represents buyer $i$’s interest in purchasing item $(i,j)$. The elements in $\mathbf{A}$ depend on the covariates and assigned treatment. Details on this sampling and computation are provided in Appendix D.1. For simplicity, we assume all buyers observe the same number of items, denoted by $M$.
Next, before sampling the outcome $Y_{ij}$, which indicates whether item $(i,j)$ is purchased, we first determine whether buyer $i$ makes any purchase at all, denoted with the binary variable $\tilde{Y}_{i}$. Conditional on a purchase occurring, we then sample which specific item the buyer purchases. We sample $\tilde{Y}_{i}$ according to the Bernoulli probability $\eta(\mathbf{A}_{i})=P(\tilde{Y}_{i}=1\mid\mathbf{A}_{i})$. The structure of the interference in this dataset is largely determined by the choice of $\eta$.
We consider three alternatives for $\eta$. The simplest is $\eta_{\text{max}}(\mathbf{A}_{i})=\max_{j}A_{ij}$ where only the most attractive item contributes the probability of a purchase by buyer $i$. We refer to $\eta_{\text{max}}$ as the max function. Next, we consider the product function $\eta_{\text{product}}(\mathbf{A}_{i})=1-\prod_{j=1}^{M}(1-A_{ij})$. This function assumes that items contribute independently to a purchase such that $P(\tilde{Y}_{i}=1)=1-P(\tilde{Y}_{i}=0)=1-\prod_{j=1}^{M}P(Y_{ij}=0)$. Lastly, the third function is inspired by position bias, commonly found in ranking systems used in e-commerce platforms (Joachims et al., 2005). We refer to this as the exponential decay function, defined as $\eta_{\text{exp-decay}}(\mathbf{A}_{i})=\sum_{j=1}^{M}(\frac{1}{2})^{\text{rank}(A_{ij})}A_{ij}$, where $\text{rank}(\cdot)$ returns the rank of the attractiveness scores for buyer $i$ in descending order. Each successive item contributes half as much as the preceding one to the probability of a purchase by buyer $i$.
For the final step, conditional on that we sample $\tilde{Y}_{i}=1$, we determine which item $(i,j)$ is purchased by sampling according to the probabilities given by the softmax function $P(Y_{ij}=1\mid\tilde{Y}_{i}=1,\mathbf{A}_{i})=e^{A_{ij}/\lambda}/\sum_{j=1}^{M}e^{A_{ij}/\lambda}$ where we set the temperature parameter $\lambda=0.1$.